nikkixmarie:

the height of a golf ball hit in the air is modeled by the equation h(t)+-16t^2+48t, where h represents the height, in feet, and t represents the number of seconds that have passed since the ball was hit. What is the height of the ball after 2 seconds.

Vocaloid:

let t = 2 and calculate h(2)

nikkixmarie:

-64

Vocaloid:

-16t^2+48t = -16(2^2) + 48(2) = ?

nikkixmarie:

t = 2,1

Vocaloid:

-16(2^2) + 48(2) = ?

nikkixmarie:

ohhh sorry i meant 32

Vocaloid:

good 32 is your answer

nikkixmarie:

thx